Repsol Q2 Pad Gas and Mud Release SITREP # 6
ALASKA DEPARTMENT OF ENVIRONMENTAL CONSERVATION
Division of Spill Prevention and Response
Prevention and Emergency Response Program
Repsol Q2 Pad Gas and Mud Release
SITREP # 6
SPILL NUMBER: 12399904601
TIME/DATE OF SPILL: Repsol reported the incident to ADEC at 10:29 AM on February 15, 2012.
TIME/DATE OF SITUATION REPORT: 2:00 PM on February 19, 2012
TYPE/AMOUNT OF PRODUCT SPILLED: Repsol estimates that approximately 42,000 gallons (1,000 barrels) of fresh-water-based drilling mud have been released to the ice drilling pad and adjacent snow-covered tundra. An unknown amount of gas has been released. The mud that was released is a mixture of minerals (primarily barium sulfate) and water. Drilling mud is used to lubricate the drill, control formation pressure, and remove cuttings from the bore. No oil has been spilled.
INCIDENT LOCATION: Qugruk #2 Pad, on the Colville River Delta, approximately 18 miles northeast of Nuiqsut and approximately 150 miles southeast of Barrow (70° 27’ 19” N, 150° 44’ 52” W).
CAUSE OF SPILL: Repsol contractor Nabors Drilling was drilling an exploratory well when the drill penetrated a shallow gas pocket at a depth of 2,523 feet, resulting in a gas kick. The gas kick drove drilling mud out of the well and through the gas diverter, which extends 75 feet from the drilling rig, onto the ice pad and adjacent snow-covered tundra. Additional mud was pumped into the well in an attempt to control it, but that mud was also forced out by the gas.
POTENTIAL RESPONSIBLE PARTY (PRP): Repsol E&P USA
RESPONSE ACTION: The well control team continues to thaw and evaluate equipment on the drilling rig. Crews are working to bypass the frozen drilling mud injection line to regain the capability to put mud into the well. They are also working to restore full function to the hydraulic system that operates the rig’s annular preventer. Today workers began clearing frozen water and drilling mud from the entrance to the drill floor, the drill floor itself and the drill cellar.
Delineation and cleanup of the off-pad spill area remain on hold pending completion of the well control work.
U.S. Fish and Wildlife Service personnel did not find evidence of a polar bear den in the spill area, but for the safety of cleanup personnel, certified bear hazers employed by Alaska Clean Seas (ACS) remain on-site round-the-clock.
Two ADEC personnel remain stationed at the Palm Pad forward command post to monitor well control work and assist with site delineation and the start of cleanup. An ADEC inspector conducted a document inspection at the rig yesterday and found that all required on-site documentation was in place.
SOURCE CONTROL: The well ceased flowing at about 9:00 PM on February 16. Preparations for a well control operation are ongoing.
RESOURCES AFFECTED: Snow-covered tundra and ice pad. Drilling mud can affect tundra plants by changing soil pH and salinity. Brackish water produced by the well may also increase soil salinity.
FUTURE PLANS AND RECOMMENDATIONS: Cleanup cannot safely begin until the well is under control. The well control team will give cleanup personnel approximately 24 hours notice in advance of completion of the well control operation. Plans are in place for delineating and cleaning up the spill, and necessary personnel crews and equipment are on standby. A reduced incident management team will continue to monitor the operation from Anchorage and two ADEC responders will remain at Palm Pad until cleanup operations are imminent, at which time additional staff will be recalled.
UNIFIED COMMAND AND PERSONNEL:
Incident Commander: Jeremy Michels, O’Brien’s Response Management
F.O.S.C.: Matt Carr, EPA
S.O.S.C.: Steve Russell, ADEC
L.O.S.C.: Gordon Brower, North Slope Borough
Field SOSC: Wes Ghormley, ADEC
WEATHER: Today: Scattered flurries, areas of fog; high temperature -6°F; east wind 20 mph; wind chill to -36°F.
Tonight: Areas of blowing snow, scattered flurries and fog; low -15°F; east wind 20 mph; wind chill to -37°F.
Monday: Scattered flurries, areas of fog; high -13°F; northeast wind to 20 mph; wind chill to -41°F.
TIME/DATE OF THE NEXT REPORT: 2:00 PM on February 20, 2012
FOR ADDITIONAL INFORMATION CONTACT: Ty Keltner, Public Information Officer, ADEC (907) 465-5009
Photographs and other spill information will be available for viewing at: http://www.state.ak.us/dec/spar/perp/index.htm
Posted: February 19, 2012